The gym is 21 years older than the auditorium. The gym is also 3 times as old as the auditorium. How old is each building?
Answer by Earlsdon(6294)

Let G = age of gym and A = age of auditorium.
G = A + 21 yrs. The gym is 21 years older than the auditorium.
G = 3A The gym is 3 times as old as the auditorium.
Since G = G, then:
A + 21 = 3A Solve for A.
2A = 21
A = 10.5 The auditorium is 10.5 years old.
G = 3A
G = 3(10.5)
G = 31.5 The gym is 31.5 years old.